# The Inadequacy of Similarity-based Privacy Metrics: Privacy Attacks against “Truly Anonymous” Synthetic Datasets

### v2

@@ 1 Introduction
Generating synthetic tabular data using machine learning algorithms has attracted growing interest from the research community [69], regulatory bodies [67, 38], government agencies [9, 92, 93], as well as investors [22, 125, 39]. The intuition is to learn the probability distribution of the real data and create new synthetic records by sampling from the trained model. This approach promises an unlimited drop-in replacement for releasing sensitive data, de-biasing, data augmentation, etc. However, models trained without robust privacy guarantees can memorize individual data points [15, 135], which enables attacks like membership and attribute inference [57, 6, 19, 115, 63, 132].
The established framework to limit information leakage from trained models is Differential Privacy (DP) [33, 35]; in the context of synthetic data, generative models should be trained while satisfying DP end to end [145, 70, 77]. However, while several synthetic data providers claim their products adhere to regulations like GDPR, HIPAA, or CCPA (see Sec. 2.2), many of them use ad-hoc heuristics to demonstrate privacy empirically rather than DP. Some combine DP with unperturbed heuristics, which breaks the end-to-end DP pipeline and ultimately negates its privacy guarantees, as our analysis will confirm. This is worrisome, as models are often trained on sensitive data in highly regulated environments such as healthcare [65]. Research papers, not only in the health domain, use models that exclusively rely on these heuristics, e.g., [103, 75, 138, 147, 98, 133, 11, 53, 113, 74, 142, 112, 71, 18, 146, 24, 144].
Most synthetic data solutions use ad-hoc heuristics based on similarity, i.e., how similar/close synthetic records are to their nearest neighbor in the train data. More precisely, if enough synthetic points are too close according to pre-configured statistical tests vs. holdout test data, either they are filtered out or the whole dataset is discarded. Otherwise, the synthetic data is considered safe. In other words, synthetic data should be similar and representative of the train data but not too close. These heuristics must be applied to every synthetic data release, as they are the only privacy mechanism some companies use.
@@
We then focus on the more complex problem of reconstructing records used to train the generative model. We introduce ReconSyn, a general black-box attack reconstructing train data from low-density regions under realistic assumptions – i.e., the adversary can only access a single fitted generative model and the metrics. ReconSyn operates by generating synthetic datasets deemed private by the metrics, and it is agnostic to the generative approach, dataset type, use case, etc. It includes two subattacks: 1) SampleAttack, which can only reconstruct data records generated by the trained model, and 2) SearchAttack, which can augment the data points obtained in the first subattack.
Experimental Evaluation. We demonstrate the effectiveness of our attacks vis-à-vis five state-of-the-art generative models for tabular data (PrivBayes [145], MST [77], DPGAN [136], PATE-GAN [70], CTGAN [137]) and five commonly used datasets (2d Gauss, Adult Small, Adult, Census, and MNIST). We show that ReconSyn reconstructs 78% to 100% of the underrepresented train data records (or outliers) with perfect precision in all settings. Some models are more vulnerable: attacking graphical models (PrivBayes, MST) requires fewer rounds (i.e., API calls to the model/metrics) to achieve similar results as GANs. We focus on outliers as they potentially correspond to the most at-risk individuals. Moreover, reconstructing outliers is inherently harder as the most difficult-to-model records likely reside in these low-density regions (as we confirm in Appendix A). Similarly, we focus on tabular data for two reasons: 1) the heuristics (and all the generative models we evaluate) are specifically designed and deployed for the tabular domain, and 2) there is an increasing number of real-world deployments in this space [9, 91, 80, 95].
@@
The next step would be to “DP-fy” the metrics while still running statistical pass/fail tests. However, as the metrics must be applied every time, this would quickly deplete the privacy budget, thus preventing unlimited data generation, one of the providers’ main selling point. Also, one could, in theory, impose a limit on the number of queries allowed for each user, but this would also prevent unlimited data generation. Moreover, synthetic data platforms are typically deployed on the client’s premises or private clouds [51, 129, 87, 59]; thus, the provider often lacks direct access to monitor or limit the number of queries. Similarly, disabling access to the metric scores would compromise the provider’s transparency and explainability, preventing users from verifying compliance claims without tangible proof of passing these tests, undermining another purported advantage of synthetic data.
@@
    We introduce ReconSyn, a novel reconstruction attack with minimal assumptions, i.e., black-box access to a single trained generative model and the privacy metrics. After applying both attack phases, SampleAttack and SearchAttack, ReconSyn reconstructs 78%–100% of the outliers in all datasets; see blue bars in Fig. 1.
@@
    We highlight the importance of moving away from reasoning about privacy in an ad-hoc, empirical way; rather, practitioners should stick to established end-to-end DP pipelines.
Figure 1: High-level overview of the performance of ReconSyn and DifferenceAttack. The privacy metrics enable privacy leakage via API calls to them (and the generative model). ReconSyn reconstructs outliers from the train data with performance varying by attack phase (SampleAttack and SearchAttack) and the number of calls. DifferenceAttack achieves a 100% success rate for membership and attribute inference ($k$ denotes the number of possible categories for the unknown attribute).
@@ 2.1 Synthetic Data and (DP) Generative Models
Synthetic Data from Generative Models. A sample dataset $D_{train}^{n}\in Z$ (consisting of $n$ iid records) is used as input to the generative model training algorithm $G(D_{train}^{n})$ during the fitting step. Next, $G(D_{train}^{n})$ updates its parameters $\theta$ to capture a representation of $P(D_{train}^{n})$ and outputs the trained model $G_{\overline{\theta}}(D_{train}^{n})$. Then, $G_{\overline{\theta}}(D_{train}^{n})$ can be used to repeatedly sample a synthetic dataset of arbitrary size $n^{\prime}$, $D_{synth}^{n^{\prime}}{\sim}G_{\overline{\theta}}(D_{train}^{n})$. Finally, we use $D_{train}^{out}\in D_{train}^{n}$ to denote the outliers or train records with low density. We refer to Appendix B for details on the datasets used in our experiments and how we define outliers for each of them.
@@
$P[{\mathcal{A}}(D)\in S]\leq\exp\left(\epsilon\right)\cdot P[{\mathcal{A}}(D^{\prime})\in S]+\delta$ (1)
@@ 2.2 Commercial Synthetic Data Solutions
Company Compliance DP Ad-hoc Heuristics
@@
Replica1 [108] SF [107]
Statice2 [117] DCR3 [116]
1Acquired by Aetion in Jan. $2$Acquired by Anonos in Nov. 2022. 3 Used for linkability and inference risk metrics [47].
Table 1: Synthetic data companies, along with whether they claim to be offering regulatory-compliant synthetic data as well as the similarity-based privacy metrics and privacy filters they use.
Over the past couple of years, we have kept track of the main synthetic data providers: Gretel, Tonic, Mostly AI, Hazy, Aindo, DataCebo, Syntegra, YData, Synthesized, Syntho, Replica Analytics, and Statice. We examined publicly available information on their claims of regulatory compliance and their approach to privacy, i.e., whether they support DP training and what heuristics they use, as summarized in Table 1. We found they mainly use five heuristics, which we describe in Sec. 2.3: three metrics, Identical Match Share (IMS), Distance to Closest Records (DCR), Nearest Neighbor Distance Ratio (NNDR), and two filters, Similarity Filter (SF) and Outlier Filter (OF).
N.B. For simplicity, we refer to the three metrics as similarity-based privacy metrics (SBPMs) and use the term (ad-hoc) heuristics to include all metrics and filters.
@@ 2.3 Similarity-based Privacy Metrics (SBPMs)
Identical Match Share (IMS) is a privacy metric that captures the proportion of identical copies between train and synthetic records. The test passes if that proportion is smaller or equal to the one between train and test datasets. IMS is used or advocated by companies [82, 118, 139, 26, 123], and scientific papers [75, 96, 4, 110, 24].
Distance to Closest Records (DCR) also compares the two sets of distances. It looks at the overall distribution of the distances to the nearest neighbors or closest records. The test passes if ($D_{train}^{n}$, $D_{synth}^{n^{\prime}}$)-5th percentile is larger or equal than the other pair. DCR is supposed to protect against settings where the train data is just slightly perturbed and presented as synthetic. Several companies [127, 82, 60, 118, 139, 116, 123], and scientific studies [103, 75, 138, 147, 96, 4, 133, 11, 71, 53, 74, 142, 112, 18, 146, 144] use it.
@@ 4 Counter-Examples
Figure 5: ReconSyn Overview. The provider 1. splits the private data into train/test, 2. fits the generative model on train data, 3. generates synthetic data (privacy filters are applied), 4. runs the privacy metrics on synthetic data. The adversary can make API calls (i.e., black-box access) to the fitted generative model and privacy metrics. They a. generate synthetic datasets, b. run them through the privacy metrics to observe the pass/fail tests and scores (if tests pass), c. reconstruct outlier train records (through SampleAttack and SearchAttack).
@@
CE3. SF & Reconstruction. We can reconstruct all outliers if Similarity Filter (SF) is applied. We assume access to an oracle possessing knowledge of the generative process. Suppose we rely on the oracle to sample 100,000 new datasets, act as if they are synthetic datasets, and use them as input to the privacy metrics. We apply SF at generation, select the datasets passing all privacy tests, and plot them in Fig. 4. Since no generative model was trained (i.e., the train data was never exposed to a model), any data directly sampled from the oracle preserves train data privacy.
However, all train outliers can immediately be detected and reconstructed from the ‘holes’ in the data—we refer to this emerging pattern as “Swiss Cheese.” This simple experiment shows that even though SF could naively be considered an additional privacy layer, filtering data out actually exposes data points. Furthermore, outliers could uniquely be identified since they are typically in low-density regions. A similar pattern could be observed if the Outlier Filter (OF) is applied. This phenomenon is also discussed in [69], although not demonstrated experimentally.
@@ 5.2 Membership and Attribute Inference Attacks
Attribute Inference. We frame attribute inference [141, 115] through the lens of multiple membership inference attacks, similar to previous work [54, 10, 56]. We assume the adversary has access to a partial target record from the train data, which we denote as $r_{t}$. In other words, they know all but one attribute of the target record and aim to infer the missing attribute. We adapt DifferenceAttack for this task as follows. The adversary makes $k$ calls, using different possible values for the unknown attribute, to the privacy metrics ($D_{synth}^{n^{\prime}}\cup r_{t_{1}}$, $D_{synth}^{n^{\prime}}\cup r_{t_{2}}$, $\ldots$, $D_{synth}^{n^{\prime}}\cup r_{t_{k}}$). The correct attribute value corresponds to the call returning the highest IMS. In principle, the attack could be extended to $t$ unknowns attributes, as done by [97]; however, the chance of multiple reconstructed records increases, potentially reducing precision.
@@ 5.3 Reconstruction Attack
ReconSyn starts by locating the regions with outliers through OutliersLocator and then launches two subattacks: 1) SampleAttack, which generates and evaluates samples drawn from the generative model, and 2) SearchAttack, which strategically examines the history of records generated in the first phase. ReconSyn’s pseudocode is provided in Algorithm 2.
@@
14:procedure SampleAttack($G_{\overline{\theta}}$, $M$, $GM$, $r_{sma}$, $n_{train}$, $c_{out}$)
15:     Init. $R_{sma}\leftarrow\emptyset$triangleright SMA recon. outliers to the empty set
16:     Init. $H_{out}\leftarrow\emptyset$triangleright history to the empty set
17:     for $r$ in $r_{sma}$ do $\triangleright$ iterate over number of rounds
18:         Generate $S\leftarrow G_{\overline{\theta}}.\text{sample}(n_{train})$triangleright new synthetic data
19:         Select $S\leftarrow\{S[i]\;|\;GM.\text{predict}(S)[i]\in c_{out}\}$
20$\triangleright$ outliers candidates
21:         Filter $S\leftarrow S\;\backslash\;H_{out}$triangleright candidates out from history
22:         Query $\text{dists}\leftarrow M(S)$triangleright metrics (augment if necessary)
23:         Update $R_{sma}\leftarrow R_{sma}\;\cup\;\{S[i]\;|\;dists[i]=0\}$
24$\triangleright$ recon. outliers
25:         Update $H_{out}\leftarrow H_{out}\;\cup\;\{\text{Zip}(S,dists)\}$triangleright history
26:     return $R_{sma},H_{out}$
27:procedure SearchAttack($M$, $GM$, $d_{sra}$, $c_{out}$, $H_{out}$)
28:     Init. $R_{sra}\leftarrow\emptyset$triangleright SRA recon. outliers to the empty set
29:     Select and sort $H^{\prime}_{out}\leftarrow\{H_{out}[i]\;|\;dists[i]\leq d_{sra}\}$
30$\triangleright$ trim history
31:     for $s,dist_{s}$ in $H^{\prime}_{out}$ do $\triangleright$ iterate over history
32:         Build $N_{s}\leftarrow\{s\;\text{with}\;s[i]\;\text{modified},\;\forall i\in[1,\text{length}(s)]\}$
33$\triangleright$ record neighboring dataset
34:         Filter $N_{s}\leftarrow N_{s}\;\backslash\;H_{out}$triangleright candidates out from history
35:         Query $\text{dists}\leftarrow M(N_{s})$triangleright metrics (augment if necessary)
36:         Update $H_{out}\leftarrow H_{out}\;\cup\;\{\text{Zip}(N_{s},dists)\}$triangleright history
37:         Select $c_{s}\leftarrow\{i\;|\;dists[i]\leq dist_{s}\}$triangleright find cols yet to be recon.
38:         for $c_{s_{i}}$ in $c_{s}$ do $\triangleright$ iterate over cols yet to be recon.
39:              Build $CC_{s}\leftarrow\{s_{i}\;|\;\forall\,val\in\text{Support}(c_{s_{i}})\}$
40$\triangleright$ col closer candidates
41:              Select $CC_{s}\leftarrow\{CC_{s}[j]\;|\;GM.\text{predict}(CC_{s})[j]\in c_{out}\}$
42$\triangleright$ outliers candidates
43:              Filter $CC_{s}\leftarrow CC_{s}\;\backslash\;H_{out}$triangleright candidates out from history
44:              Query $\text{dists}\leftarrow M(CC_{s})$triangleright metrics (augment if necessary)
45:              Update $R_{sra}\leftarrow R_{sra}\;\cup\;\{CC_{s}[j]\;|\;dists[j]=0\}$
46$\triangleright$ recon. outliers
47:              Update $H_{out}\leftarrow H_{out}\;\cup\;\{\text{Zip}(CC_{s},dists)\}$triangleright history
48:     return $R_{sra}$
@@
Table 2: Comparison between DifferenceAttack and state-of-the-art membership inference attacks on a target train outlier on Adult Small: AUC score, number of API calls, and runtime.
@@
Table 3: Overview of ReconSyn’s performance (SampleAttack and SearchAttack) with different datasets and against different models (all DP models are trained with $\epsilon=\infty$). NB: the ‘1 call’ column refers to SampleAttack with only one API call to the metrics.

### v3

@@ 1 Introduction
Generating synthetic tabular data using machine learning algorithms has attracted growing interest from the research community [67], regulatory bodies [65, 37], government agencies [9, 88, 89], as well as investors [21, 120]. The intuition is to learn the probability distribution of the real data and create new synthetic records by sampling from the trained model. This approach promises an unlimited drop-in replacement for releasing sensitive data, de-biasing, data augmentation, etc. However, models trained without robust privacy guarantees can memorize individual data points [15, 130], which enables attacks like membership and attribute inference [55, 6, 18, 110, 61, 127].
The established framework to limit information leakage from trained models is Differential Privacy (DP) [32, 34]; in the context of synthetic data, generative models should be trained while satisfying DP end to end [139, 68, 75]. However, while several synthetic data providers claim their products adhere to regulations like GDPR, HIPAA, or CCPA (see Sec. 2.2), many of them use ad-hoc heuristics to demonstrate privacy empirically rather than DP. Some combine DP with unperturbed heuristics, which breaks the end-to-end DP pipeline and ultimately negates its privacy guarantees, as our analysis confirms. This is worrisome, as these products are typically trained on sensitive data and are already deployed in highly regulated environments, e.g., healthcare [63].
Many synthetic data solutions use ad-hoc heuristics based on similarity, i.e., how similar/close synthetic records are to their nearest neighbor in the train data. More precisely, if enough synthetic points are too close according to pre-configured statistical tests vs. holdout test data, either they are filtered out or the whole dataset is discarded. Otherwise, the synthetic data is considered safe. In other words, synthetic data should be similar and representative of the train data but not too close. These heuristics must be applied to every synthetic data release, as they are the only privacy mechanism some companies use.
@@
We then focus on the more complex problem of reconstructing records used to train the generative model. We introduce ReconSyn, a black-box attack reconstructing train data from low-density regions under realistic assumptions – i.e., the adversary can only access a single fitted generative model and the metrics. ReconSyn operates by generating synthetic datasets deemed private by the metrics, and it is agnostic to the generative approach, dataset type, use case, etc. It includes two subattacks: 1) SampleAttack, which can only passively reconstruct data records generated by the trained model, and 2) SearchAttack, which can actively augment the data points obtained in the first subattack.
Experimental Evaluation. We demonstrate the effectiveness of our attacks vis-à-vis five popular generative models for tabular data (PrivBayes [139], MST [75], DPGAN [131], PATE-GAN [68], CTGAN [132]) and five commonly used datasets (2d Gauss, Adult Small, Adult, Census, and MNIST). We show that ReconSyn reconstructs 78% to 100% of the underrepresented train data records (or outliers) with perfect precision in all settings, primarily due to the privacy leakage from the metrics. Some models are more vulnerable: attacking graphical models (PrivBayes, MST) requires fewer rounds (i.e., API calls to the model/metrics) to achieve similar results as GANs. We focus on outliers as they potentially correspond to the most at-risk individuals. Moreover, reconstructing outliers is inherently harder as the most difficult-to-model records likely reside in these low-density regions (as we confirm in Appendix A). Similarly, we focus on tabular data, as: 1) the heuristics (and all studied generative models) are specifically designed and deployed for the tabular domain, and 2) there is an increasing number of real-world deployments in this space [9, 87, 91].
@@
The next step would be to ‘DP-fy’ the metrics while still running statistical pass/fail tests. However, as the metrics must be applied every time, this would quickly deplete the privacy budget, thus preventing unlimited data generation, one of the providers’ main selling point. Also, one could, in theory, impose a limit on the number of queries allowed for each user, but this would also prevent unlimited data generation. Moreover, synthetic data platforms are typically deployed on the client’s premises or private clouds [49, 124, 83, 57]; thus, the provider often lacks direct access to monitor or limit the number of queries. Similarly, disabling access to the metric scores would compromise the provider’s transparency and explainability, preventing users from verifying compliance claims without tangible proof.
@@
    We introduce ReconSyn, a novel reconstruction attack with minimal assumptions, i.e., black-box access to a single trained generative model and the privacy metrics. After applying both attack phases, SampleAttack and SearchAttack, ReconSyn reconstructs 78%–100% of the outliers in all datasets; see the blue bars in Fig. 1.
@@
    We stress the need to move away from reasoning about privacy in an ad-hoc, empirical way; rather, practitioners should adopt established end-to-end DP pipelines.
Figure 1: Overview of the reconstruction success rate of ReconSyn and DifferenceAttack. ReconSyn reconstructs outliers from the train data with success varying by attack phase (SampleAttack and SearchAttack) and the number of calls. DifferenceAttack achieves a 100% success for membership and attribute inference ($k$ denotes the number of possible categories for the unknown attribute).
@@ 2.1 Synthetic Data and (DP) Generative Models
Synthetic Data from Generative Models. A sample dataset $D_{train}^{n}$ (consisting of $n$ iid records) is used as input to the generative model training algorithm $G(D_{train}^{n})$ during the fitting step. Next, $G(D_{train}^{n})$ updates its parameters $\theta$ to capture a representation of $P(D_{train}^{n})$ and outputs the trained model $G_{\overline{\theta}}(D_{train}^{n})$. Then, $G_{\overline{\theta}}(D_{train}^{n})$ can be used to repeatedly sample a synthetic dataset of arbitrary size $n^{\prime}$, $D_{synth}^{n^{\prime}}{\sim}G_{\overline{\theta}}(D_{train}^{n})$. Finally, we use $D_{train}^{out}\in D_{train}^{n}$ to denote the outliers or train records with low density. We refer to Appendix B for details on the datasets used in our experiments and how we define outliers for each of them.
@@
$\small P[{\mathcal{A}}(D)\in S]\leq\exp\left(\epsilon\right)\cdot P[{\mathcal{A}}(D^{\prime})\in S]+\delta$ (1)
@@ 2.2 Commercial Synthetic Data Solutions
Company Claimed DP Ad-hoc
Compliance Models Privacy Heuristics
@@
Replica2 [103] SF [102]
Statice3 [112] DCR4 [111]
1Acquired by SAS in Nov. 2024.  2Acquired by Aetion in Jan. 2022. 3Acquired by Anonos in Nov. 2022. 4Used for linkability and inference risk metrics [45].
Table 1: Synthetic data companies and whether they claim to be offering regulatory-compliant synthetic data or DP models, as well as the similarity-based privacy metrics/filters they use.
We have been tracking the main synthetic data providers (Gretel, Tonic, Mostly AI, Hazy, Aindo, DataCebo, Syntegra, YData, Synthesized, Syntho, Replica Analytics, and Statice) and examined publicly available information on their claims of regulatory compliance and their approach to privacy, i.e., whether they support DP training and what heuristics they use. As summarized in Table 1, they mainly use five heuristics, which we describe in Sec. 2.3: three metrics, Identical Match Share (IMS), Distance to Closest Records (DCR), Nearest Neighbor Distance Ratio (NNDR), and two filters, Similarity Filter (SF) and Outlier Filter (OF). N.B. For simplicity, we refer to the three metrics as similarity-based privacy metrics (SBPMs) and use the term (ad-hoc) heuristics to include all metrics and filters.
@@
Note that, besides commercial products, peer-reviewed research outputs have also used models that exclusively rely on the heuristics – see, e.g., [98, 73, 141, 93, 128, 11, 51, 108, 72, 136, 107, 69, 140, 23, 138].
@@ 2.3 Similarity-based Privacy Metrics (SBPMs)
Identical Match Share (IMS) is a privacy metric that captures the proportion of identical copies between train and synthetic records. The test passes if that proportion is smaller or equal to the one between train and test datasets. IMS is used or advocated by companies [79, 113, 133, 25, 118], and scientific papers/blogposts [73, 4, 105, 23].
Distance to Closest Records (DCR) also compares the two sets of distances. It looks at the overall distribution of the distances to the nearest neighbors or closest records. The test passes if ($D_{train}^{n}$, $D_{synth}^{n^{\prime}}$)-5th percentile is larger or equal than the other pair. DCR is supposed to protect against settings where the train data is just slightly perturbed and presented as synthetic. Several companies [122, 79, 58, 113, 133, 111, 118], and scientific studies/blogposts [98, 73, 141, 4, 128, 11, 69, 51, 72, 136, 107, 140, 138] use it.
@@ 4 Counter-Examples
Figure 5: ReconSyn Overview. The provider 1. splits the private data into train/test, 2. fits the generative model on train data, 3. generates synthetic data (privacy filters are applied), 4. runs the privacy metrics on synthetic data. The adversary can make API calls (i.e., black-box access) to the fitted generative model and privacy metrics. They a. generate synthetic datasets, b. run them through the privacy metrics to observe the pass/fail tests and scores (if tests pass), c. reconstruct train data outliers (through SampleAttack and SearchAttack).
@@
CE3. SF & Reconstruction. We can reconstruct all outliers if Similarity Filter (SF) is applied. We assume access to an oracle possessing knowledge of the generative process. Suppose we use the oracle to sample 100,000 synthetic datasets, apply SF, and select the datasets passing all privacy tests (we plot them in Fig. 4). Since no generative model was trained (i.e., the train data was never exposed to a model), any data directly sampled from the oracle preserves privacy.
However, all train outliers can immediately be detected and reconstructed from the ‘holes’ in the data—we refer to this emerging pattern as “Swiss Cheese.” This simple experiment shows that even though SF could naively be considered an additional privacy layer, filtering data out could actually expose data points. Furthermore, outliers could uniquely be identified since they are typically in low-density regions. A similar pattern could be observed if the Outlier Filter (OF) is applied. This phenomenon is also discussed in [67], although not demonstrated experimentally. In Appendix D, we also discuss how this simple experiment fails to scale to higher dimensions unless specific records are targeted.
@@ 5.2 Membership and Attribute Inference Attacks
Attribute Inference. We frame attribute inference [135, 110] through the lens of multiple membership inference attacks, similar to previous work [52, 10, 54]. We assume the adversary has access to a partial target record from the train data, which we denote as $r_{t}$. In other words, they know all but one attribute of the target record and aim to infer the missing attribute. We adapt DifferenceAttack for this task: the adversary makes $k$ calls, using different possible values for the unknown attribute, to the privacy metrics ($D_{synth}^{n^{\prime}}\cup r_{t_{1}}$, $D_{synth}^{n^{\prime}}\cup r_{t_{2}}$, $\ldots$, $D_{synth}^{n^{\prime}}\cup r_{t_{k}}$). The correct attribute value corresponds to the call with highest IMS. In principle, the attack could be extended to $t$ unknown attributes, as in [92]; however, the chance of multiple reconstructed records increases, potentially reducing precision.
@@ 5.3 Reconstruction Attack
ReconSyn starts by locating the regions with outliers through OutliersLocator and then launches two subattacks: 1) SampleAttack, which passively generates and evaluates samples drawn from the generative model, and 2) SearchAttack, which actively and strategically examines the history of records generated in the first phase. ReconSyn’s pseudocode is provided in Algorithm 2.
@@
14:end procedure
15:procedure SampleAttack($G_{\overline{\theta}}$, $M$, $GM$, $r_{sma}$, $n_{train}$, $c_{out}$)
16:     Init. $R_{sma}\leftarrow\emptyset$triangleright SMA recon. outliers to the empty set
17:     Init. $H_{out}\leftarrow\emptyset$triangleright history to the empty set
18:     for $r$ in $r_{sma}$ do $\triangleright$ iterate over number of rounds
19:         Generate $S\leftarrow G_{\overline{\theta}}.\text{sample}(n_{train})$triangleright new synthetic data
20:         Select $S\leftarrow\{S[i]\;|\;GM.\text{predict}(S)[i]\in c_{out}\}$
21$\triangleright$ outliers candidates
22:         Filter $S\leftarrow S\;\backslash\;H_{out}$triangleright candidates out from history
23:         Query $\text{dists}\leftarrow M(S)$triangleright metrics (augment if necessary)
24:         Update $R_{sma}\leftarrow R_{sma}\;\cup\;\{S[i]\;|\;dists[i]=0\}$
25$\triangleright$ recon. outliers
26:         Update $H_{out}\leftarrow H_{out}\;\cup\;\{\text{Zip}(S,dists)\}$triangleright history
27:     end for
28:     return $R_{sma},H_{out}$
29:end procedure
30:procedure SearchAttack($M$, $GM$, $d_{sra}$, $c_{out}$, $H_{out}$)
31:     Init. $R_{sra}\leftarrow\emptyset$triangleright SRA recon. outliers to the empty set
32:     Select and sort $H^{\prime}_{out}\leftarrow\{H_{out}[i]\;|\;dists[i]\leq d_{sra}\}$
33$\triangleright$ trim history
34:     for $s,dist_{s}$ in $H^{\prime}_{out}$ do $\triangleright$ iterate over history
35:         Build $N_{s}\leftarrow\{s\;\text{with}\;s[i]\;\text{modified},\;\forall i\in[1,\text{length}(s)]\}$
36$\triangleright$ record neighboring dataset
37:         Filter $N_{s}\leftarrow N_{s}\;\backslash\;H_{out}$triangleright candidates out from history
38:         Query $\text{dists}\leftarrow M(N_{s})$triangleright metrics (augment if necessary)
39:         Update $H_{out}\leftarrow H_{out}\;\cup\;\{\text{Zip}(N_{s},dists)\}$triangleright history
40:         Select $c_{s}\leftarrow\{i\;|\;dists[i]\leq dist_{s}\}$triangleright find cols yet to be recon.
41:         for $c_{s_{i}}$ in $c_{s}$ do $\triangleright$ iterate over cols yet to be recon.
42:              Build $CC_{s}\leftarrow\{s_{i}\;|\;\forall\,val\in\text{Support}(c_{s_{i}})\}$
43$\triangleright$ col closer candidates
44:              Select $CC_{s}\leftarrow\{CC_{s}[j]\;|\;GM.\text{predict}(CC_{s})[j]\in c_{out}\}$
45$\triangleright$ outliers candidates
46:              Filter $CC_{s}\leftarrow CC_{s}\;\backslash\;H_{out}$triangleright candidates out from history
47:              Query $\text{dists}\leftarrow M(CC_{s})$triangleright metrics (augment if necessary)
48:              Update $R_{sra}\leftarrow R_{sra}\;\cup\;\{CC_{s}[j]\;|\;dists[j]=0\}$
49$\triangleright$ recon. outliers
50:              Update $H_{out}\leftarrow H_{out}\;\cup\;\{\text{Zip}(CC_{s},dists)\}$triangleright history
51:         end for
52:     end for
53:     return $R_{sra}$
54:end procedure
@@
Table 2: Comparison between DifferenceAttack and state-of-the-art membership inference attacks on a target train outlier, Adult Small: AUC score, number of API calls, and runtime.
@@
Table 3: Overview of ReconSyn’s reconstruction success rate (SampleAttack and SearchAttack) with different datasets and against different models (all DP models are trained with $\epsilon=\infty$). NB: ‘1 call’ refers to SampleAttack with only one API call to the metrics.